# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ A.1 List of Hypotheses
[…]
@@
[…]
@@
      W[SAT]-SETH: For all $\varepsilon>0$ there is no algorithm which takes as input a (general) Boolean formula $\phi$ of size $s$ on $n$ variables and an integer $k$ and decides if $\phi$ is satisfiable in time $n^{(1-\varepsilon)k}s^{O(1)}$.
@@ A.2 Relations between Classes of Hypotheses
We apply the following standard trick exhaustively: as long as $\phi$ contains a clause of arity $r>3$, say $C=(\ell_{1}\lor\ell_{2}\lor\ldots\lor\ell_{r})$, introduce $r-3$ new variables $z_{1},\ldots,z_{r-3}$ and replace $C$ with the clauses $(\ell_{1}\lor\ell_{2}\lor z_{1})\land(\neg z_{1}\lor\ell_{3}\lor z_{2})\land(\neg z_{2}\lor\ell_{4}\lor z_{3})\land\ldots\land(\neg z_{r-3}\lor\ell_{r-1}\lor\ell_{r})$. Observe now that the variables of $C$ form a clique in the primal graph, while the new variables induce a path whose neighborhood is contained in this clique. We can therefore apply the claim, with $\mathcal{K}$ being the set of all clauses on which we applied this transformation.
We note here two further easy observations justifying the organization of hypotheses we have made in Figure 1.

### v2

@@ A.1 List of Hypotheses
      $\log\!\textrm{td}$M-SETH: The same as the previous statement but for tree-depth at most $c\log|\phi|$.
@@
      $\log\!\textrm{pw}$M-SETH: The same as the previous statement but for pathwidth at most $c\log|\phi|$.
@@
      W[SAT]-SETH: For all $\varepsilon>0$ there is no algorithm which takes as input a (general) Boolean formula $\phi$ of size $s$ on $n$ variables and an integer $k$ and decides if $\phi$ is satisfiable by an assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}s^{O(1)}$.
@@ A.2 Relations between Classes of Hypotheses
We apply the following standard trick exhaustively: as long as $\phi$ contains a clause of arity $r>3$, say $C=(\ell_{1}\lor\ell_{2}\lor\ldots\lor\ell_{r})$, introduce r3r-3 new variables $z_{1},\ldots,z_{r-3}$ and replace $C$ with the clauses $(\ell_{1}\lor\ell_{2}\lor z_{1})\land(\neg z_{1}\lor\ell_{3}\lor z_{2})\land(\neg z_{2}\lor\ell_{4}\lor z_{3})\land\ldots\land(\neg z_{r-3}\lor\ell_{r-1}\lor\ell_{r})$. Observe now that the variables of $C$ form a clique in the primal graph, while the new variables induce a path whose neighborhood is contained in this clique. We can therefore apply the claim, with $\mathcal{K}$ being the set of all clauses on which we applied this transformation.
We note here two further easy observations justifying the organization of hypotheses we have made in Figure˜1.